In the following circuit, V = 55 volts, R1 = 3 Ohms, R2 = 5 Ohms, R3 = 3 Ohms, R4 = 10 Ohms, and R5 = 6 Ohms. What is the current through R 3 in amps?

Ri w 3 R2 + 4 mm R5
Answer #1

R,=32 ww 55 V - 1 R5 bri Apply KuL in loop 0 -vrir,+(91-92) R2 + 9, R5=0 (Rit R₂ +R5)9,- R₂ iz =v . (3+5+6)91-592=55 149,-512